Circular No. 6956

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cr IN PGC 57979 T. Puckett and A. Langoussis, Mountain Town, GA, report the dis-

covery of an apparent supernova (mag 17.6) on an unfiltered CCD frame taken with the Puckett Observatory 0.30-m automated supernova patrol telescope on June 28.09 UT. SN 1998cr is located at α = 16^{h}23^{m}17^{s}.25, δ = +37^{o}12′47′′.5 (equinox 2000.0), which is 0′′.5 west and 7′′.5 south of the nucleus of PGC 57979. The object was also present on CCD frames taken on June 29.10 at mag 17.5. The object was not present on previous CCD frames taken on June 24.2 (limiting mag 18.75). The object does not appear on the Palomar Sky Survey scans. M. Richmond, Rochester Institute of Technology, reports that SN 1998cr was present on a 300-s R-band KAIT image taken on June 29.275 UT.

SUPERNOVAE 1998ce, 1998ch, 1998ck M. Phillips, Cerro Tololo Interamerican Observatory (CTIO), reports

that he obtained spectroscopic typing of three supernovae with the CTIO 4-m Blanco telescope. SN 1998ce (cf. IAUC 6922) is a type-II event; its spectrum, obtained on June 27.05 UT, closely resembles that of SN 1992H, 50 days after explosion (see Clocchiatti et al. 1996, A.J. 111, 1286). SN 1998ch (cf. IAUC 6917) is also of type II, and appears to be approximately the same age as SN 1998ce; the spectrum of SN 1998ch was obtained on June 27.07, and a spectrum of its host galaxy shows Hα emission at a redshift z ∼ 0.047, consistent with membership in Abell 3559. The spectrum of SN 1998ck (cf. IAUC 6925), obtained on June 27.03, shows this object to be a type-Ia supernova, ∼ 18 days after maximum; a spectrum of its host galaxy, ESO 434-G20, yields z ∼ 0.038.

SUPERNOVA 1998S IN NGC 3877 Additional CCD photometry by D. Hanžl, Brno, Czech Republic, (cf.

IAUC 6885, 6924): June 2.904 UT, V = 15.15 ± 0.05; 17.918, V = 15.33 ± 0.05, B − V = +0.81 ± 0.05; 19.918, 15.45 ± 0.03, +0.7 ± 0.2; 20.982, 15.40 ± 0.08, –; 21.893, 15.43 ± 0.03, –; 25.889, 15.38 ± 0.06, +0.62 ± 0.1.

SGR 1627–41

Corrigendum. On IAUC 6950, line 5, for May read June

1998 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green